The table shows the marginal cost C(x), the marginal revenue R(x) for producing x items. The third column, P(x), is partiaP(x) -21 O 43 NOT 64 10 43 43 40 16 43 70 43 43 90208143 43 - 165
Answer #1

given 43 OOOOX x I cw pliky | plx, l 64 lo 142 143 107 43 27 7043430 go 208 43 1-65 16 9) (6 pix) = RX) - c(X) when x=10 -